OpenStudy (anonymous):

solve for x 8^x-8=7^4x

OpenStudy (anonymous):

this is a transcendental eq....can't solve it directly....use approximation or numerical methods.

OpenStudy (anonymous):

logarithm?

OpenStudy (anonymous):

sorry i have to solve it in log form

OpenStudy (anonymous):

you could try but theat would not change the transcencdental nature of the equation

OpenStudy (anonymous):

is left hand side 8 to the power (x-8) or (8 t0 power x) - 8

OpenStudy (anonymous):

Write the exact answer using base-10 logarithms. or no solution

OpenStudy (anonymous):

Is it 8^(x-8)? or 8^x -8

OpenStudy (anonymous):

you'd get an x inside the log i.e. as an argument

OpenStudy (anonymous):

they are both the same

OpenStudy (anonymous):

no...they're not

OpenStudy (anonymous):

They aren't.

OpenStudy (anonymous):

certianly not same

OpenStudy (anonymous):

It would make a big difference if it was 8 to the power x-3

OpenStudy (anonymous):

rip:use the equation editor for clarity...

OpenStudy (anonymous):

8 to the power of x-3

OpenStudy (anonymous):

then 7 to the power 4x

OpenStudy (anonymous):

sorry guys =(

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

\[8^{x-3} = 7^{4x}\] ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

x = 9log(2)/(3log(2) - log(7))

OpenStudy (anonymous):

note you get that by taking a log on both sides

OpenStudy (anonymous):

sorry....x = 9log(2)/(3log(2) - 4log(7))

OpenStudy (anonymous):

so thats it?

OpenStudy (anonymous):

sorry for the earlier trascendental analysis .. I thought that the last x was multiplied by 7^4

OpenStudy (anonymous):

yep...hope it was of some help.

OpenStudy (anonymous):

thanks!!!!

hero (hero):

\[x = -\frac{8 \log{8}}{4 \log{7} - \log{8}} = -2.91637\]

OpenStudy (anonymous):

(x-3) ln 8 = 4x ln 7 x-3 / 4x = ln7 / ln 8= 0.93578 x = -1.094

OpenStudy (anonymous):

Slight mistake: it's \[3 \log{8} /(\log 8 - 4\log 7)\] You switched the signs on -4log 7 and log8

OpenStudy (anonymous):

Also, this does nothing for her understanding. Here's a more detailed explanation of how to solve this problem: 1) \[8^(x-3)=7^4x\] 2) Insert log functions and carry exponents to left of log expression --> \[(x-3) \log8 = 4x \log 7\] 3) Distribute the terms in the exponents to the log expressions --> \[x \log8 - 3 \log8 = 4x \log7\] 4. Combine like-terms and factor -->\[x(\log8 - 4 \log7) = 3 \log8\] 5. Isolate the variable x --> \[x = 3 \log8/(\log8 - 4 \log 7)\] 6. Solve with calculator... unless you can evaluate logarithmic values in without one... Answer: -1.09363726

hero (hero):

I didn't get it wrong. The original post was 8^(x-8) = 7^(4x)

hero (hero):

I wasn't aware that it was changed to three instead of 8

OpenStudy (anonymous):

i dont quite see what u mean - i'll plug the results into the original equations

hero (hero):

Jimmy Rep, you don't see what the original equation was? It's posted right at the top

OpenStudy (anonymous):

I misread. The answer really is -2.916366027 Method is correct, though. Sorry